Solve by the elimination method.
3x+4y=3
6x+8y=6

Answer by stanbon(75887)

Solve by the elimination method.
3x+4y=3
6x+8y=6
----------------------
Multiply thru the 1st equation by 2 to get:
6x + 8y = 6
6x + 8y = 6
-----------------
Subtract to get:
0 = 0
----
Although the original equation look different,
one is a multiple of the other.
---
The solution is all points of the form (x,(-3/4)x+(3/4))
where x is any Real Number.
==========================================================
Cheers,
Stan H.
===============